# Phonon-mediated spin-polarized superconductivity in altermagnets

### v1

@@ Acknowledgements.
$H^{\prime}=e^{-\zeta S}He^{\zeta S}=H_{0}+\zeta\left(H_{1}+\left[H_{0},S\right]\right)+\zeta^{2}\left[H_{1},S\right]+\frac{1}{2}\zeta^{2}\left[\left[H_{0},S\right],S\right]+\mathcal{O}\left(\zeta^{3}\right).$ (S12)
@@
$H^{\prime}=H_{0}+\frac{1}{2}\zeta^{2}[H_{1},S]+\mathcal{O}\left(\zeta^{3}\right).$ (S14)
To find the S satisfying (S13), we start by looking at $H_{0}$ and $H_{1}$,
$H_{0}$ $=\sum_{\textbf{k}\sigma}\epsilon_{\textbf{k}\sigma}d_{\textbf{k}\sigma}^{\dagger}d_{\textbf{k}\sigma}+\sum_{\textbf{q}\lambda}\hbar\omega_{\textbf{q}\lambda}\left(a_{\textbf{q}\lambda}^{\dagger}a_{\textbf{q}\lambda}+\frac{1}{2}\right),$ (S15)
$H_{1}$ $=\sum_{\textbf{k}\textbf{q}\sigma\lambda}g_{\lambda\sigma}\left(\textbf{k},\textbf{k}+\textbf{q}\right)\left(a_{-\textbf{q},\lambda}^{\dagger}+a_{\textbf{q}\lambda}\right)d_{\textbf{k}+\textbf{q},\sigma}^{\dagger}d_{\textbf{k}\sigma}.$ (S16)
@@
$S=\sum_{\textbf{k}\textbf{q}\sigma\lambda}g_{\lambda\sigma}\left(\textbf{k},\textbf{k}+\textbf{q}\right)\left(x_{\textbf{k}\textbf{q}\sigma\lambda}a_{-\textbf{q},\lambda}^{\dagger}+y_{\textbf{k}\textbf{q}\sigma\lambda}a_{\textbf{q}\lambda}\right)d_{\textbf{k}+\textbf{q},\sigma}^{\dagger}d_{\textbf{k}\sigma},$ (S17)
solving for the coefficients $x_{\textbf{k}\textbf{q}\sigma\lambda}$ and $y_{\textbf{k}\textbf{q}\sigma\lambda}$ by applying the condition (S13) gives
$x_{\textbf{k}\textbf{q}\sigma\lambda}$ $=\frac{1}{\epsilon_{\textbf{k}\sigma}-\epsilon_{\textbf{k}+\textbf{q},\sigma}-\hbar\omega_{\textbf{q}\lambda}},$ (S18)
$y_{\textbf{k}\textbf{q}\sigma\lambda}$ $=\frac{1}{\epsilon_{\textbf{k}\sigma}-\epsilon_{\textbf{k}+\textbf{q},\sigma}+\hbar\omega_{\textbf{q}\lambda}}.$ (S19)
Note that since we have assumed the scattering to be contained on the FS, the two electron energies in the denominators will cancel. We can now find the effective electron-electron interaction in (S14), where we only keep terms corresponding to electron-electron interactions and neglect higher-order phonon terms,
$H_{\mathrm{eff}}=\frac{1}{2}\sum_{\textbf{k}\textbf{q}\sigma}\sum_{\lambda\textbf{k}^{\prime}\sigma^{\prime}}g_{\lambda\sigma}\left(\textbf{k},\textbf{k}+\textbf{q}\right)g_{\lambda\sigma^{\prime}}\left(\textbf{k}^{\prime},\textbf{k}^{\prime}-\textbf{q}\right)\left(y_{\textbf{k}\textbf{q}\sigma\lambda}-x_{\textbf{k}\textbf{q}\sigma\lambda}\right)d_{\textbf{k}+\textbf{q},\sigma}^{\dagger}d_{\textbf{k}^{\prime}-\textbf{q},\sigma^{\prime}}^{\dagger}d_{\textbf{k}\sigma}d_{\textbf{k}^{\prime}\sigma^{\prime}}.$ (S20)
As explained in the main text, we consider spin-polarized, zero-momentum Cooper pairs, which set the conditions $\sigma^{\prime}=\sigma$ and $^{\prime}=-\textbf{k}$ k. With these assumptions, the interaction takes the form
$H_{\mathrm{eff}}=\sum_{\textbf{k}\textbf{k}^{\prime}\sigma}V_{\textbf{k}^{\prime}\textbf{k}\sigma}d_{\textbf{k}^{\prime}\sigma}^{\dagger}d_{-\textbf{k}^{\prime},\sigma}^{\dagger}d_{-\textbf{k},\sigma}d_{\textbf{k}\sigma},$ (S21)
@@
$V_{\textbf{k}^{\prime}\textbf{k}\sigma}=-\sum_{\lambda}\frac{g_{\sigma\lambda}(\textbf{k},\textbf{k}^{\prime})g_{\sigma\lambda}(-\textbf{k},-\textbf{k}^{\prime})}{\hbar\omega_{\textbf{k}^{\prime}-\textbf{k},\lambda}},$ (S22)
@@
$H_{\mathrm{eff}}$ $=\frac{1}{2}\sum_{\textbf{k}\textbf{k}^{\prime}\sigma}\bar{V}_{\textbf{k}^{\prime}\textbf{k}\sigma}d_{\textbf{k}^{\prime}\sigma}^{\dagger}d_{-\textbf{k}^{\prime},\sigma}^{\dagger}d_{-\textbf{k},\sigma}d_{\textbf{k}\sigma},$ (S23)
$2\bar{V}_{\textbf{k}^{\prime}\textbf{k}\sigma}$ $=V_{\textbf{k}^{\prime}\textbf{k}\sigma}+V_{-\textbf{k}^{\prime},-\textbf{k},\sigma}-V_{-\textbf{k}^{\prime},\textbf{k}\sigma}-V_{\textbf{k}^{\prime},-\textbf{k},\sigma}.$
S.XI Details of the effective interaction on the Fermi surface
In this section, we discuss in more detail the momentum dependence of the effective electron-electron interaction on the FS. We introduce the quantity $\Lambda_{k_{\|}k^{\prime}_{\|}\sigma}$ via the linearized Fermi-surface averaged gap-equation [39]
@@
Figure S1: The first column shows the matrix elements $\Lambda_{k_{\|}k^{\prime}_{\|}\sigma}$ for spin-down, with k given by the yellow dashed line, corresponding to the midpoint between A and B in the respective plot in the third column, and $\textbf{k}^{\prime}$ is given by the position along the path. The second column has k as marked in the respective third column plots, corresponding to the maximum of the gap. The third column shows the gap for spin-down. The rows are structured such that in the first row (a-c), we have used $\left(t_{2}=0,\mu=0.1t\right)$, in the second (d-f) $\left(t_{2}=0,\mu=0.25t\right)$, in the third (g-i) $\left(t_{2}\approx 0.21t,\mu=0.1t\right)$, and in the fourth (j-l) $\left(t_{2}\approx 0.21t,\mu=0.25t\right)$. N indicates the momentum with the most negative value for the gap. Table SI lists the remaining parameters.
In the first row of Fig. S1 with $t_{2}=0$ and $\mu=0.1t$, Fig. 4 (a) in the main text shows that the dimensionless coupling is $\lambda_{m}\approx 0.01$, which is too low to give a non-negligible $T_{c}$. From (S24) and Fig. S1(a), we see that $\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}$ has a disadvantageous sign, as discussed in the main text. It is also most substantial for $^{\prime}=\pm\textbf{k}$ k along the $\bm{\Gamma}\textbf{X}$ line. Hence, it is natural that the eigenvector with the largest eigenvalue has zero gap on the $\bm{\Gamma}\textbf{X}$ line. Figure S1(a) shows $\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}$ when k is placed at the maximal gap amplitude. For that value of k, we see that $|\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}|$ is not maximal at $^{\prime}=\pm\textbf{k}$ k, rather it is maximal for $\textbf{k}^{\prime}$ on the $\bm{\Gamma}\textbf{X}$ line.
By comparing the coupling in Fig. S1(b) with the gap in Fig. S1(c), we see how appropriate sign changes of the gap generate an eigenvector with a positive eigenvalue $\lambda_{m}$. From Fig. S1(b), we see that $|\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}|$ with k placed in the maximum positive gap ($\Delta_{\textbf{k}\downarrow}>0$) and $\textbf{k}^{\prime}$ placed in the maximum negative gap ($\Delta_{\textbf{k}^{\prime}\downarrow}<0$, $\textbf{k}^{\prime}=\mathrm{N}$), is smaller than the maximum $|\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}|$ in Fig. S1(a). As a result, the eigenvector is barely able to satisfy the requirement of positive eigenvalue, and $\lambda_{m}$ is small.
Then, in the second row of Fig. S1 with $t_{2}=0$ and $\mu=0.25t$, Fig. 4 (a) of the main text shows that $\lambda_{m}\approx 0.20$, which is larger than the first row by a considerable amount, leading to a non-vanishing critical temperature. This may be understood from the fact that the maximal magnitude of $|\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}|$ is much greater in Fig. S1(e) than in Fig. S1(d). Also, unlike the first row [Fig. S1(b)], the peak of $|\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}|$ in Fig. S1(e) occurs close to where k is placed in the maximum positive gap and $\textbf{k}^{\prime}$ is placed in the maximum negative gap. Hence, the sign change of the gap function takes significant advantage of the coupling, giving a greater eigenvalue $\lambda_{m}$.
The situation in the third and fourth rows with $t_{2}\approx 0.21t$, is very similar to the second row. In the third row, however, the difference in magnitudes $|\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}|$ between column one and two is about the same as in the first row, but an important difference is that the values of $|\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}|$ are larger, making it possible to have a higher $\lambda_{m}$. More importantly, the shape of $\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}$ in Fig. S1(h) is more similar to Figs. S1(e) and S1(k) than Fig. S1(b). This gives a markedly different momentum structure of the gap in rows two, three, and four [Figs. S1(f), S1(i), and S1(l)] compared to row one [Fig. S1(c)]. This difference in the momentum dependence of the coupling and gap function enables a larger eigenvalue $\lambda_{m}$.
The increased $\lambda_{m}$ when including NNN hopping is understood from the effect of non-zero $t_{2}$ on the inverse slope of the FS, shown in Fig. 2 (b) of the main text. Compared to $t_{2}=0$, the slope is not as strongly peaked along the $\bm{\Gamma}\textbf{X}$ line, reducing the adverse effects of $|\Lambda_{k_{\|}k^{\prime}_{\|}\downarrow}|$ with $^{\prime}=\pm\textbf{k}$ k on the $\bm{\Gamma}\textbf{X}$ line. As mentioned in the main text, a non-zero $t_{2}$ also enables more electron-phonon processes which increases the coupling strength.
S.XII Parameters
@@
Table SI shows the value of parameters that are the same throughout the letter. Parameters that are varied, are given in the caption of their respective figures.
Table SI: List of parameters

### v2

@@ C.1 Effective electron-electron interaction
$H^{\prime}$ $=e^{-\zeta S}He^{\zeta S}$ (22)
$=H_{0}+\zeta\left(H_{1}+\left[H_{0},S\right]\right)$
$+\zeta^{2}\left[H_{1},S\right]+\frac{1}{2}\zeta^{2}\left[\left[H_{0},S\right],S\right]+\mathcal{O}\left(\zeta^{3}\right).$
@@
$H^{\prime}=H_{0}+\frac{1}{2}\zeta^{2}[H_{1},S]+\mathcal{O}\left(\zeta^{3}\right).$ (24)
To find the S satisfying (23), we start by looking at $H_{0}$ and $H_{1}$,
$H_{0}$ $=\sum_{\textbf{k}\sigma}\varepsilon_{\textbf{k}\sigma}d_{\textbf{k}\sigma}^{\dagger}d_{\textbf{k}\sigma}+\sum_{\textbf{q}\lambda}~\hbar\omega_{\textbf{q}\lambda}\left(a_{\textbf{q}\lambda}^{\dagger}a_{\textbf{q}\lambda}+\frac{1}{2}\right),$ (25)
$H_{1}$ $=\sum_{\textbf{k}\textbf{q}\sigma\lambda}g_{\lambda\sigma}\left(\textbf{k},\textbf{k}+\textbf{q}\right)\left(a_{-\textbf{q},\lambda}^{\dagger}+a_{\textbf{q}\lambda}\right)d_{\textbf{k}+\textbf{q},\sigma}^{\dagger}d_{\textbf{k}\sigma}.$ (26)
@@
$S=$ $\sum_{\textbf{k}\textbf{q}\sigma\lambda}g_{\lambda\sigma}\left(\textbf{k},\textbf{k}+\textbf{q}\right)$ (27)
$\times$ $\left(x_{\textbf{k}\textbf{q}\sigma\lambda}a_{-\textbf{q},\lambda}^{\dagger}+y_{\textbf{k}\textbf{q}\sigma\lambda}a_{\textbf{q}\lambda}\right)d_{\textbf{k}+\textbf{q},\sigma}^{\dagger}d_{\textbf{k}\sigma},$
solving for the coefficients $x_{\textbf{k}\textbf{q}\sigma\lambda}$ and $y_{\textbf{k}\textbf{q}\sigma\lambda}$ by applying the condition (23) gives
$x_{\textbf{k}\textbf{q}\sigma\lambda}$ $=\frac{1}{\varepsilon_{\textbf{k}\sigma}-\varepsilon_{\textbf{k}+\textbf{q},\sigma}-\hbar\omega_{\textbf{q}\lambda}},$ (28)
$y_{\textbf{k}\textbf{q}\sigma\lambda}$ $=\frac{1}{\varepsilon_{\textbf{k}\sigma}-\varepsilon_{\textbf{k}+\textbf{q},\sigma}+\hbar\omega_{\textbf{q}\lambda}}.$ (29)
Note that since we have assumed the scattering to be contained on the FS, the two electron energies in the denominators will cancel. We can now find the effective electron-electron interaction in (24), where we only keep terms corresponding to electron-electron interactions and neglect higher-order phonon terms,
$H_{\mathrm{eff}}=$ $\frac{1}{2}\sum_{\textbf{k}\textbf{q}\sigma}\sum_{\lambda\textbf{k}^{\prime}\sigma^{\prime}}g_{\lambda\sigma}\left(\textbf{k},\textbf{k}+\textbf{q}\right)g_{\lambda\sigma^{\prime}}\left(\textbf{k}^{\prime},\textbf{k}^{\prime}-\textbf{q}\right)$ (30)
$\times\left(y_{\textbf{k}\textbf{q}\sigma\lambda}-x_{\textbf{k}\textbf{q}\sigma\lambda}\right)d_{\textbf{k}+\textbf{q},\sigma}^{\dagger}d_{\textbf{k}^{\prime}-\textbf{q},\sigma^{\prime}}^{\dagger}d_{\textbf{k}\sigma}d_{\textbf{k}^{\prime}\sigma^{\prime}}.$
As explained in Sec. V, we consider spin-polarized, zero-momentum Cooper pairs, which set the conditions $\sigma^{\prime}=\sigma$ and $\textbf{k}^{\prime}=-\textbf{k}$. With these assumptions, the interaction takes the form
$H_{\mathrm{eff}}=\sum_{\textbf{k}\textbf{k}^{\prime}\sigma}V_{\textbf{k}^{\prime}\textbf{k}\sigma}d_{\textbf{k}^{\prime}\sigma}^{\dagger}d_{-\textbf{k}^{\prime},\sigma}^{\dagger}d_{-\textbf{k},\sigma}d_{\textbf{k}\sigma},$ (31)
@@
$V_{\textbf{k}^{\prime}\textbf{k}\sigma}=-\sum_{\lambda}\frac{g_{\sigma\lambda}(\textbf{k},\textbf{k}^{\prime})g_{\sigma\lambda}(-\textbf{k},-\textbf{k}^{\prime})}{\hbar\omega_{\textbf{k}^{\prime}-\textbf{k},\lambda}},$ (32)
@@
$H_{\mathrm{eff}}$ $=\frac{1}{2}\sum_{\textbf{k}\textbf{k}^{\prime}\sigma}\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}^{\prime}\textbf{k}\sigma}d_{\textbf{k}^{\prime}\sigma}^{\dagger}d_{-\textbf{k}^{\prime},\sigma}^{\dagger}d_{-\textbf{k},\sigma}d_{\textbf{k}\sigma},$ (33)
$2\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}^{\prime}\textbf{k}\sigma}$ $=V_{\textbf{k}^{\prime}\textbf{k}\sigma}+V_{-\textbf{k}^{\prime},-\textbf{k},\sigma}-V_{-\textbf{k}^{\prime},\textbf{k}\sigma}-V_{\textbf{k}^{\prime},-\textbf{k},\sigma}.$
Appendix D Details of the effective interaction on the Fermi surface
In this section, we discuss in more detail the momentum dependence of the effective electron-electron interaction on the FS. We introduce the quantity $\Lambda_{k_{\|}k^{\prime}_{\|}\sigma}$ via the linearized Fermi-surface averaged gap equation [39, 58, 83]
@@ Appendix D Details of the effective interaction on the Fermi surface
Figure 5: The first column shows the matrix elements $\Lambda_{k_{\|}k^{\prime}_{\|}\sigma}$ for spin-down, with k given by the yellow dashed line, corresponding to the midpoint between A and B in the respective plot in the third column, and $\textbf{k}^{\prime}$ is given by the position along the path. The second column has k as marked in the respective third column plots, corresponding to the maximum of the gap. The third column shows the gap for spin-down. The rows are structured such that in the first row (a-c), we have used $\left(t_{2}=0,\mu=0.1t\right)$, in the second (d-f) $\left(t_{2}=0,\mu=0.25t\right)$, in the third (g-i) $\left(t_{2}\approx 0.21t,\mu=0.1t\right)$, and in the fourth (j-l) $\left(t_{2}\approx 0.21t,\mu=0.25t\right)$. N indicates the momentum with the most negative value for the gap. Table 1 lists the remaining parameters.
[…]
Appendix E Finite-momentum Cooper pairs
Figure 6: The box marks the BZ, solid lines denote states with energies within a phonon bandwidth of the Fermi level, and the dashed lines indicate which FS is shifted. The overlap between the shifted FS and the opposite spin-FS is marked in purple. The spin-up FS is shifted in (a) and (c), while for (b) and (d), it is spin-down. The vector starting in the $\Gamma$-point is the finite-momentum vector, showing how the FS is shifted. The finite-momentum vector is chosen to maximize the overlap in (a) and (b), while for (c) and (d), it sets the overlap at the ends of the ellipses highlighted by the purple arrow. In all the cases, $\mu=0.1$ and $t_{2}=0$, as these values give relatively large overlap. See Table 1 for the remaining parameters.
In the Sec. V, we have focused on spin-polarized pairing. However, there is still a possibility for both singlet and spin-zero triplet states. To form different-spin Cooper pairs, a finite-momentum state is required, effectively shifting one of the spin-FS in Fig. 2(a) on top of the other, as illustrated in Fig. 6. Even in the case of maximum overlap, there is a significantly smaller area as opposed to the spin-polarized case, where the whole region between the solid lines can form Cooper pairs. In the cases of Fig. 6 (a) and (b), where we have maximized the overlap, the intersection area is $0.073$ times the area of the spin-polarized case. While for the case of Fig. 6 (c) and (d), the intersection area is $0.016$ times the spin-polarized area. Let us compare this to Fig. 1(b) in Ref. [54]. In their model, the ellipses have a rather large overlap for the case comparable to our Fig. 6 (c) and (d). The Lieb lattice model has much sharper edges of the ellipses here, and much more dispersive bands in this region compared to the rest of the ellipse, both of which significantly reduces the likelihood of finite-momentum pairing compared to the model of Ref. [54] and other models that have studied the competition of zero-momentum and finite-momentum pairing in altermagnets [55]. Also note that the case in Fig. 6 (a) and (b) is comparable to Fig. 1(a) in Ref. [54] with 4 discrete overlap points of the FSs, only that in our case it also requires finite momentum since the spin up and spin down FSs have no overlap originally. Such a state is not expected to be competitive at strong spin splitting [54], which is the case we focus on.
Appendix F Numerical parameters
@@
Table 1 shows the value of parameters that are the same throughout the paper. Parameters varied are given in the caption of their respective figures. The number of sampling points varies with the FS length. We have a uniform spacing of $6\times 10^{-3}/a$ between neighboring points. The value in Tab. 1 is a typical number.
Table 1: List of parameters